OpenStudy (blackops2luvr):

which property is shown? - 7 + 24 = 24 + (-7)

OpenStudy (skullpatrol):

$$\huge a+b=b+a$$

OpenStudy (blackops2luvr):

\(\Huge \color{blue}{\text{Practice~time!~<3}}\)

OpenStudy (blackops2luvr):

....

OpenStudy (blackops2luvr):

\(\Huge \color{orange}{\text{....}}\)

OpenStudy (blackops2luvr):

C

OpenStudy (skullpatrol):

The order in which you add two numbers does not affect the result.

OpenStudy (blackops2luvr):

\(\Huge \color{orange}{\text{Commutative?}}\)

OpenStudy (blackops2luvr):

What is the value of the expression x + (–y) when x = –3.8 and y = 1.7? A. –5.5 B. –2.1 C. 2.1 D. 5.5